OpenStudy (anonymous):

Solve -x+4y=33 x+5y=57

OpenStudy (lgbasallote):

system... -x + 4y = 33 x+ 5y = 57 9y = 90 y = 10 x = 7

Directrix (directrix):

-x+4y=33 x+5y=57 --------- 0 + 9y = 90 y = 10 x + 5y = 57 x + 5(10) = 57 x = 7